A bicycle is traveling north at 5.0 m/s. The mass of the wheel, 2.0 kg, is uniformly distributed along the rim, which has a radius of 20 cm. What are the magnitude and direction of the angular momentum of the wheel about its axle?
1 : 2.0 kg • m2/s towards the west
2 : 5.0 kg • m2/s vertically upwards
3 : 2.0 kg • m2/s towards the east
4 : 5.0 kg • m2/s towards the east
5 : 5.0 kg • m2/s towards the west
1
Expert's answer
2020-06-01T14:20:40-0400

A bicycle is travelling north at 5 m/s. The mass of the wheel, 2 kg, is uniformly distributed along the rim, which has a radius of 20 cm.

moment of inertia of the wheel ="mr^2=2\times(\frac{20}{100})^2=\frac{2}{25}kg\ m^2"



"v=\omega r\\5=\omega \times0.2\\\omega =25 rad/s"


The angular momentum is, L=Iω = "\frac2{25}\times 25=2kg\ m^2/sec"